# MLLM4PUE: Toward Universal Embeddings in Computational Pathology through Multimodal LLMs

### v1

@@ Abstract
Pathology plays a critical role in diagnosing a wide range of diseases, yet existing approaches often rely heavily on task-specific models trained on extensive, well-labeled datasets. These methods face sustainability challenges due to the diversity of pathologies and the labor-intensive nature of data collection. To address these limitations, we highlight the need for universal multimodal embeddings that can support multiple downstream tasks. Previous approaches often involve fine-tuning CLIP-based models, which handle images and text separately, limiting their ability to capture complex multimodal relationships. Additionally, these models are evaluated across diverse datasets without a unified benchmark for assessing multimodal embeddings in pathology. To address these challenges, we propose MLLM4PUE, a novel framework that leverages Multimodal Large Language Models (MLLMs) to generate Pathology Universal Embeddings. The MLLM4PUE framework not only facilitates robust integration of images and text but also enhances understanding and fusion capabilities across various tasks. We further introduce the Pathology Multimodal Embedding Benchmark (PMEB), a comprehensive benchmark designed to assess the quality of pathology multimodal embeddings. PMEB comprises 15 original tasks drawn from 14 datasets, organized into three meta-tasks: retrieval, classification, and composed retrieval. Experimental results demonstrate the superiority of MLLM4PUE, illustrating MLLM-based models can effectively support a wide range of downstream tasks and unify the research direction for foundation models in pathology.
@@ 1 Introduction
Figure 1: Performance comparison across benchmark datasets, showing that our proposed MLLM4PUE achieves state-of-the-art (SOTA) results across all tasks. Weight F1 (wF1) is reported for classification task. For the retrieval and composed retrieval task, we report recall@5 scores. This highlights the advanced multimodal fusion capabilities of MLLM4PUE.
@@
Pathology remains the gold standard for diagnosing a wide range of diseases [42, 32, 8]. With rapid advancements in artificial intelligence (AI), computational pathology has emerged as a promising field, using AI to address challenges across various tasks such as cancer diagnosis [29, 39], metastasis detection [44, 22], mutation prediction [37, 9, 33], and survival analysis [41, 40, 45]. These achievements, however, rely on task-specific models trained on large, well-labeled datasets. Given the hundreds of tumor types cataloged in the WHO classification111tumourclassification.iarc.who.int/ and the labor-intensive nature of data collection and annotation [12], developing separate, high-performing models for each pathology task is impractical, especially for rare diseases with limited data [24, 30]. Furthermore, while many efforts focus on image data, pathology often involves natural language, including diagnostic reports and research publications [6, 7, 31, 14]. Thus, there is a growing need to shift toward task-agnostic models that generate universal multimodal embeddings capable of supporting a broad range of pathology tasks.
@@
In natural image processing, recent studies such as CLIP [34] have demonstrated remarkable success by aligning visual and language embeddings through contrastive learning, achieving strong performance across diverse downstream tasks. Building on these advances, several studies [30, 16, 38, 17, 31] have adapted CLIP-based models for pathology, training on extensive image-text pairs from public sources to produce multimodal embeddings that enable zero-shot learning across various pathology tasks. However, these methods encounter two significant challenges. First, CLIP [34] processes images and text independently, limiting its ability to capture complex interactions between modalities [19]. Second, evaluations of these models have mostly centered on isolated tasks, such as classification or retrieval [30, 38, 17], which separate image and text components and do not fully assess multimodal integration capabilities. Moreover, inconsistencies in datasets and sample selection criteria across studies lead to non-reproducible and incomparable results, emphasizing the need for a more integrated approach and a standardized benchmarking framework to advance multimodal embeddings in pathology.
To tackle the first challenge, we propose MLLM4PUE, a framework that uses a Multimodal Large Language Model (MLLM) backbone to generate universal multimodal embeddings for pathology. To the best of our knowledge, this is the first use of an MLLM backbone in this field, representing a significant advancement over previous CLIP-based methods. Unlike CLIP [34], which processes images and text separately, MLLM4PUE employs a transformer-based architecture to fully integrate these modalities, allowing it to learn from complex image-text relationships. This fusion enhances the model’s adaptability and effectiveness across a range of pathology tasks, making it particularly suited for generalizing to diverse pathology scenarios [28, 27].
@@
  • We introduce MLLM4PUE, the first framework utilizing MLLMs to generate universal multimodal embeddings in pathology. This framework integrates multimodal information within a unified model, converting any combination of images and text into robust embeddings.
@@
  • MLLM4PUE effectively represents multimodal information, consistently surpassing the performance of all baseline models across all evaluated tasks, as demonstrated by the quantitative results in Fig. 1.
  • Our approach offers new insights into computational pathology by adopting MLLMs for universal embeddings. Future research can build on this work to unify and streamline the development of models for diverse pathology tasks.
2 Related work
2.1 Multimodal embedding learning
Creating effective multimodal embeddings remains a challenging area of research. The CLIP model [34] addresses this by employing separate encoders for images and text, aligning them in a common space through contrastive learning. This approach has influenced many subsequent models, such as BLIP [26] and CoCa [43]. While these models excel in a variety of tasks, the use of distinct encoders restricts their capacity to fully integrate visual and textual data, which is a key requirement for tasks that involve combined visual-language inputs, such as search queries involving both images and text.
With the rise of large language models (LLMs), Multimodal Large Language Models (MLLMs) have extended LLMs to process multiple data modalities, achieving notable progress in understanding and reasoning across diverse input types [28, 27, 25]. Although MLLMs exhibit strong capabilities in interpreting multimodal content and following complex instructions, there is still limited research on their application in creating effective multimodal embeddings. The recent E5-V [19] model fine-tunes MLLMs with text-based data to improve embedding quality. Building on this, we have adapted the approach for pathology by using domain-specific image-text data, resulting in substantial performance improvements.
2.2 Multimodal embeddings for computational pathology
In computational pathology, recent studies have adopted the CLIP [34] architecture and training methodology, leveraging paired image-text data within a contrastive learning framework to align similar image-text embeddings while separating dissimilar ones [30, 16, 38, 17, 18]. For instance, PLIP [16] fine-tunes CLIP on large collections of image-text pairs sourced from platforms like Twitter and other public datasets. Similarly, PathCLIP [38] and QuiltNet [17] scale up pathology-specific data for fine-tuning, while CONCH [31] employs over 1.17 million image-caption pairs, making it the largest dataset used for training to date.
Despite these advancements, these methods fine-tune either CLIP [34] or CoCa [43] using pathology image-caption pairs and rely on separate encoders for processing images and text. In contrast, we propose an MLLM-based approach to capture universal multimodal embeddings for pathology by integrating image and text data within a unified model. This integrated framework enhances performance across a wide range of computational pathology tasks, offering a more holistic and effective solution.
3 Methodology
3.1 Pathology multimodal embedding enchmark
@@
CRC-100k [20] I T Colorectal cancer 100,000 1,500
@@
Kathercolon [21] I T Colorectal cancer tissue 7,180 1,500
LC25000 [4] I T Colon adenocarcinoma 10,000 1,500
LC25000 [4] I T Lung adenocarcinoma 15,000 1,500
@@
Renalcell [5] I T Renal tissue 36,687 1,500
Sicap [36] I T Gleason pattern 12,081 1,500
Skincancer [23] I T Skin cancer tissue 129,369 1,500
Wsss4luad [15] I T LUAD tissue 4,693 1,500
@@
We propose PMEB (Pathology Multimodal Embedding Benchmark), a comprehensive benchmark designed to evaluate multimodal embeddings in pathology across a diverse range of tasks. These tasks include cancer tissue classification [2, 4, 23], Gleason pattern grading [36], and cell identification [10], among other tasks. PMEB comprises 15 original tasks from 14 datasets, organized into three meta-tasks: retrieval, classification, and composed retrieval. All tasks are reformulated to assess multimodal embeddings, where the model receives an instructional prompt and a query, which may consist of text, images, or both, and selects the correct target from a set of options.
@@
In the retrieval task, the query can be either an image or text, with the target being the corresponding paired text or image. For the classification task, queries are images with targets as text labels representing various classes. The composed retrieval task involves queries that combine an image with text formatted as a question, with the target being the expected answer. This composed retrieval task, specifically designed for multimodal embedding evaluation, incorporates integrated visual and language inputs and has not been applied in previous methods. Detailed dataset statistics are provided in Table  1.
@@
Due to the large size of certain classification datasets, we have standardized their sizes. For datasets containing more than 1,500 samples, we employ the CONCH model [31] to calculate similarity scores between each image and its corresponding class label text. We then select the top 500 images with the highest scores, the bottom 500 with the lowest scores, and a random sample of 500 from the remaining data, ensuring the original category distribution is preserved. For the composed retrieval task, we select only open-ended questions from the data source, where answers are more complex than a simple “yes” or “no”. Additionally, since there is an overlap between questions and answers in the VQA data source, we utilize ChatGPT [1] for data cleaning to ensure clarity and accuracy in task evaluation. Further details and examples from each dataset within PMEB are provided in Supplementary Materials.
@@ 3.2 Multimodal Embeddings with MLLM
We first introduce how our proposed MLLM4PUE leverages MLLMs for universal multimodal embedding learning. Unlike CLIP, which directly generates embeddings for multimodal inputs, MLLMs are inherently generative models and require a different approach to produce embeddings. Inspired by E5-V [19], our solution uses a prompt-based strategy specifically designed to guide MLLMs in embedding multimodal data.
@@ 3.4 Zero-shot Transfer Evaluation
For zero-shot classification, we adopt a prompt-based method inspired by the CLIP model [34]. In this approach, each class name is expanded into a sentence using a specific template. For instance, the class name “Colon adenocarcinoma” is converted into the sentence “An H&E image of Colon adenocarcinoma” using the template “An H&E image of { }.” We apply this method to create sentences for all class names. Our model then computes embeddings for these sentences and the test images, extracting the last token of the MLLM output as described in Section 3.2. The similarity between these embeddings is calculated as outlined in Section 3.3, and test image labels are assigned based on the highest similarity scores. To account for variability in templates, we randomly select one from the 21 templates in CONCH [31] and repeat this process 100 times for each dataset. The templates are exhibited in Supplementary Materials.
@@
Table 2: Performance comparison with baseline methods on two retrieval datasets, reporting recall metrics at different thresholds. Task i2t and t2i denote image-to-text and text-to-image retrieval, respectively. MLLM4PUE-O, MLLM4PUE-P, and MLLM4PUE-Q denote our model pre-trained on Openpath [16], PathCap [38], Quilt1M [17], respectively. Bold values indicate the best performance.
@@ 4 Experiments
Figure 4: Comparison of zero-shot patch classification performance across different models, measured by weighted F1 metric. Each box represents 100 data points, with each point indicating the performance on one template. The upper, middle, and lower lines of each box denote the first, median, and third quartiles, respectively. MLLM4PUE-O, MLLM4PUE-P, and MLLM4PUE-Q denote our model pre-trained on Openpath [16], PathCap [38], Quilt1M [17], respectively.
4.1 Training Datasets
We leverage three public datasets, Openpath [16], PathCap [38] and Quilt1M [17], to construct a robust training foundation for pathology image-text pairs.
The Openpath dataset [16] is assembled by gathering data from Twitter and LAION using popular pathology hashtags. Since PLIP[16] provided only Twitter IDs, we retrieve the corresponding data from Twitter and LAION using these IDs, resulting in a dataset of 138,874 image-text pairs after denoising. The PatchCap dataset [38] comprises 207K high-quality samples from authoritative sources, and we use all provided image-text samples. Similarly, the Quilt1M dataset [17] aggregates pathology image-text pairs from several public sources. As many images in Quilt1M are associated with multiple captions, we concatenate the captions for each image to create comprehensive entries. Through these careful preparations, we compile an extensive training dataset consisting of 593,838 image-text pairs.
4.2 Baselines
We evaluate the performance of our proposed MLLM4PUE, against several recent SOTA approaches across various downstream tasks. Specifically, we compare with three currently developed CLIP-based methods: PLIP [16], PathCLIP [38], and QuiltNet [17]. We also compare our method with E5-V [19], a contemporary model that leverages vision-language models for multimodal embedding tasks. It proposes a single-modality training approach, where the model is trained exclusively on text pairs.
4.3 Implementation Details and Metrics
We use LLaVA-NeXT [25] as our backbone MLLM. To save the GPU memory, we use QLoRA [11] and gradient checkpointing with DeepSpeed ZeRO-2, and all the images are resized to 336 × 336 pixels. The temperature is set to 0.02. All experiments are run on six H100 GPUs.

### v2

# Toward Universal Embeddings in Digital Pathology through Multimodal LLMs

@@ Abstract
Pathology plays a critical role in diagnosing a wide range of diseases, yet existing approaches often rely heavily on task-specific models trained on extensive, well-labeled datasets. These methods face sustainability challenges due to the diversity of pathologies and the labor-intensive nature of data collection. To address these limitations, we highlight the need for universal multimodal embeddings that can support multiple downstream tasks. Previous approaches involve fine-tuning CLIP-based models, which handle images and texts separately, limiting their ability to capture complex multimodal relationships. Additionally, these models are evaluated across diverse datasets without a unified benchmark. In this paper, we explore the possibility of applying Multimodal Large Language Models (MLLMs) to generate pathology universal embeddings to address these challenges. Our contributions can be summarized in the following aspects: 1) We propose MLLM4PUE, a novel framework that leverages MLLMs to generate embeddings for various pathology downstream tasks. 2) We further introduce the Pathology Multimodal Embedding Benchmark (PMEB), a comprehensive benchmark designed to assess the quality of pathology multimodal embeddings, which comprises 16 original tasks drawn from 15 datasets. 3) Extensive experimental results demonstrate the superiority of MLLM4PUE, illustrating MLLM-based models can effectively support a wide range of downstream tasks and unify the research direction for foundation models in pathology.
@@ 1 Introduction
Figure 1: Performance comparison across benchmark datasets, showing that our proposed MLLM4PUE achieves baseline methods across most of tasks. Weight F1 (wF1) is reported for classification task. For the retrieval and composed retrieval task, we report recall@5 scores. This highlights the advanced multimodal fusion capabilities of MLLM4PUE.
@@
Pathology remains the gold standard for diagnosing a wide range of diseases [41, 32]. With advancements in artificial intelligence (AI), digital pathology has emerged as a promising field, using AI to address challenges across various tasks such as cancer diagnosis [29, 38], metastasis detection [43, 21], mutation prediction [8, 33], and survival analysis [40, 39, 44]. These achievements, however, rely on models trained on large, well-labeled datasets. Given the hundreds of tumor types cataloged in the WHO classification111tumourclassification.iarc.who.int/ and the labor-intensive nature of data collection and annotation [11], developing separate models for each pathology task is impractical, especially for rare diseases with limited data [23, 30]. Furthermore, while many efforts focus on image, pathology often involves text, including diagnostic reports and research papers [6, 7, 31, 13]. Thus, there is a growing need to shift toward task-agnostic models that generate universal multimodal embeddings capable of supporting different pathology tasks.
@@
Due to the success of CLIP [34] model, several studies [30, 15, 37, 16, 31] have adapted CLIP-based models for pathology, training on extensive image-text pairs to extract embeddings that enable zero-shot evaluation across various pathology tasks. However, these methods encounter two significant challenges. First, CLIP [34] processes images and text independently, limiting its ability to capture complex interactions between modalities [18]. Additionally, CLIP [34] can only handle text inputs with a maximum length of 74 tokens, which restricts its ability to leverage longer and more detailed textual descriptions. Second, evaluations of these models have mostly centered on isolated tasks, such as classification or retrieval [30, 37, 16], which separate image and text components and do not fully assess multimodal integration capabilities. Moreover, inconsistencies in datasets and sample selection criteria across studies lead to non-reproducible and incomparable results. These challenges emphasize the need for a more integrated model and a standardized benchmarking framework to advance multimodal embeddings in pathology.
To tackle the first challenge, we propose MLLM4PUE, a framework that uses a MLLM to generate universal multimodal embeddings for pathology. Unlike CLIP [34], which processes images and text separately, MLLM employs a transformer-based architecture to fully integrate these modalities, allowing it to learn from complex image-text relationships [28, 27]. Moreover, MLLM can handle longer textual inputs compared to CLIP, overcoming the 74-token limitation and enabling richer contextual understanding, which is important in pathology area. To harness the capability of MLLM in generating embeddings, we take the following steps. First, we employ summarization prompts to guide the MLLM to output embeddings. Next, we fine-tune MLLM with a LoRA module using pathology image-text pairs, adapting it specifically for the pathology domain.
@@
  • We introduce MLLM4PUE, a framework converting pathology images and texts into robust embeddings within a unified model. This is the first time to explore the possibility of re-purposing MLLM for extracting embeddings in pathology area.
@@
  • MLLM4PUE effectively represents multimodal information, consistently surpassing the performance of baseline models across amounts of evaluated tasks, as demonstrated by the quantitative results in Fig. 1.
  • Our approach offers new insights into digital pathology by adopting MLLM for universal embeddings. Future research can build on this work to unify the development of models for diverse pathology tasks.
@@
CRC-100k [19] I T Colorectal cancer 100,000 1,000
Databiox [19] I T Colorectal cancer 922 922
@@
Kathercolon [20] I T Colorectal cancer tissue 7,180 1,000
LC25000 [4] I T Colon adenocarcinoma 10,000 1,000
LC25000 [4] I T Lung adenocarcinoma 15,000 1,000
@@
Renalcell [5] I T Renal tissue 36,687 1,000
Sicap [36] I T Gleason pattern 12,081 1,000
Skincancer [22] I T Skin cancer tissue 129,369 1,000
Wsss4luad [14] I T LUAD tissue 10,091 1,000
@@
2 Related Work
Vision-language Models in Digital Pathology. In digital pathology, recent studies have adopted the CLIP [34] model, leveraging paired image-text data within a contrastive learning framework to align similar image-text embeddings while separating dissimilar ones [30, 15, 37, 16, 17]. For instance, PLIP [15] fine-tunes CLIP on large collections of image-text pairs sourced from platforms like Twitter and other public datasets. Similarly, PathCLIP [37] and QuiltNet [16] scale up pathology-specific data for training. CONCH [31] employs over 1.17 million image-caption pairs to fine-tune CoCa [42] model. Despite these advancements, these methods fine-tune either CLIP [34] or CoCa [42] using pathology image-caption pairs and rely on separate encoders for processing images and text. In contrast, we propose an MLLM-based framework to capture universal multimodal embeddings for pathology by integrating image and text data within a unified model. This integrated framework enhances performance across a wide range of digital pathology tasks, offering a more holistic and effective solution.
Multimodal Embedding Learning. Creating effective multimodal embeddings remains a challenging area of research. CLIP  [34] addresses this by employing separate encoders for images and text, aligning them in a common space through contrastive learning. This approach has influenced many subsequent models, such as BLIP [25] and CoCa [42]. While these models excel in a variety of tasks, the use of distinct encoders restricts their capacity to fully integrate visual and textual data, which is a key requirement for tasks that involve combined visual-language inputs, such as search queries involving both images and text.
With the rise of large language models (LLMs), MLLMs have extended LLMs to process multiple data modalities, achieving notable progress in understanding and reasoning across diverse input types [28, 27, 24]. Although MLLMs exhibit strong capabilities in interpreting multimodal content and following complex instructions, there is still limited research on their application in creating embeddings. The recent work E5-V [18] fine-tunes an LLM with summarization prompts and text-only data to extract embeddings, then integrates a vision module to obtain multimodal embeddings for zero-shot multimodal retrieval tasks. However, for pathological adaptation, E5-V [18] requires a domain-specific LLM pretrained on pathology instruction data, which demands detailed and standardized annotations. In contrast, our framework, MLLM4PUE can fine-tune MLLM with a LoRA module using existing pathology image-text pairs, bypassing these prohibitive requirements.
3 Methodology
3.1 Pathology Multimodal Embedding Benchmark
@@
We propose PMEB, a comprehensive benchmark designed to evaluate pathology multimodal embeddings. PMEB include cancer tissue classification [2, 4, 22], Gleason pattern grading [36], and cell identification [9], among other tasks. As shown in Table 1, PMEB comprises 15 original tasks from 14 datasets, organized into three meta-tasks: retrieval, classification, and composed retrieval. All tasks are reformulated to assess multimodal embeddings, where the model receives an instructional prompt and a query, which may consist of text, images, or both, and selects the correct target from a set of options.
@@
In the retrieval task, the query can be either an image or text, with the target being the corresponding paired text or image. For the classification task, queries are images with targets as text labels representing various classes. The composed retrieval task involves queries that combine an image with text formatted as a question, with the target being the expected answer. This composed retrieval task, specifically designed for multimodal embedding evaluation, incorporates integrated visual and language inputs and has not been applied in previous methods.
@@
Due to the large size of some datasets, we standardize their sizes. For datasets containing more than 1,000 samples (except Osteo [3]), we choose a random sample of 1,000, ensuring the original category distribution is preserved. For the composed retrieval task, we select only open-ended questions from the data source, where answers are more complex than a simple “yes” or “no”. Additionally, since there is an overlap between questions and answers in the VQA data source, we utilize ChatGPT [1] for data cleaning to ensure clarity and accuracy in task evaluation. Further details of each dataset within PMEB are provided in Supplementary Materials.
@@ 3.2 Multimodal Embeddings with MLLM
We first introduce how our proposed MLLM4PUE leverages MLLMs for universal multimodal embedding learning. Unlike CLIP, which directly generates embeddings for multimodal inputs, MLLMs are inherently generative models and require a different approach to produce embeddings. We uses a prompt-based strategy specifically designed to guide MLLMs in embedding multimodal data.
@@ 3.4 Zero-shot Transfer Evaluation
For zero-shot classification, we adopt a prompt-based method inspired by the CLIP model [34]. In this approach, each class name is expanded into a sentence using a specific template. For instance, the class name “Colon adenocarcinoma” is converted into the sentence “An H&E image of Colon adenocarcinoma” using the template “An H&E image of { }.” We apply this method to create sentences for all class names. Our model then computes embeddings for these sentences and the test images, extracting the last token of the MLLM output as described in Section 3.2. The similarity between these embeddings is calculated as outlined in Section 3.3, and test image labels are assigned based on the highest similarity scores.
@@
Table 2: Performance comparison with baseline methods on two retrieval datasets, reporting recall metrics at different thresholds. Task i2t and t2i denote image-to-text and text-to-image retrieval, respectively. MLLM4PUE-O, MLLM4PUE-P, and MLLM4PUE-Q denote our model pre-trained on Openpath [15], PathCap [37], Quilt1M [16], respectively. Bold values indicate the best performance. Same color represents models trained on the same dataset.
@@ 4 Experiments
Training datasets. We leverage three public datasets, Openpath [15], PathCap [37] and Quilt1M [16], to construct a robust training foundation for pathology image-text pairs. The Openpath dataset [15] is assembled by gathering data from Twitter and LAION using popular pathology hashtags. Since PLIP[15] provided only Twitter IDs, we retrieve the corresponding data from Twitter and LAION using these IDs, resulting in a dataset of 138,874 image-text pairs after denoising. The PatchCap dataset [37] comprises 207K high-quality samples from authoritative sources, and we use all provided image-text samples. Similarly, the Quilt1M dataset [16] aggregates pathology image-text pairs from several public sources. As many images in Quilt1M are associated with multiple captions, we concatenate the captions for each image to create comprehensive entries. Through these careful preparations, we compile an extensive training dataset consisting of 593,838 image-text pairs.
Baseline Methods and Metrics. We evaluate the performance of our proposed MLLM4PUE against several recent baseline models across various downstream tasks. Specifically, we compare it with three pathology CLIP-based models: PLIP [15], PathCLIP [37], and QuiltNet [16], all of which are trained on the same dataset as our method. Additionally, we compare our approach with E5-V [18], a method that fine-tunes an LLM with summarization prompts and text-only data to extract embeddings, then integrates a vision module to obtain multimodal embeddings for zero-shot multimodal retrieval tasks. For zero-shot classification tasks, we also compare our method with CONCH [31], a model trained on over 1.17 million private image-caption pairs, whereas our method is trained on fewer than 0.6 million pairs.
@@
Implementation Details. Our framework is implemented in Pytorch and we use LLaVA-NeXT-8B [24] as our backbone MLLM. To save the GPU memory, we use QLoRA [10] and gradient checkpointing with DeepSpeed ZeRO-2 to fine-tune the MLLM, and all the images are resized to 336 × 336 pixels. The temperature is set to 0.02. All experiments are run on six H100 GPUs with a gradient accumulation batch size of 144 and a learning rate of 4e-4.
Model Bach
Data
biox
Digest
Digest
CRC
100k
Kather
7k
LC
colon
LC
lung
Osteo
Renal
cell
Sicap
Skin
cancer
Wsss
4luad
E5-V [18] 0.100 0.162 0.197 0.049 0.224 0.206 0.333 0.209 0.300 0.051 0.230 0.017 0.132
CONCH [31] 0.511 0.413 0.835 0.824 0.680 0.625 0.890 0.650 0.382 0.393 0.359 0.593 0.772
PLIP [15] 0.124 0.194 0.696 0.252 0.503 0.504 0.697 0.743 0.554 0.413 0.080 0.311 0.394
MLLM4PUE-O 0.561 0.437 0.784 0.480 0.642 0.509 0.958 0.744 0.581 0.507 0.284 0.411 0.636
PathCLIP [37] 0.528 0.282 0.785 0.707 0.611 0.492 0.937 0.887 0.668 0.359 0.267 0.332 0.648
MLLM4PUE-P 0.577 0.388 0.799 0.900 0.763 0.783 0.937 0.886 0.751 0.543 0.502 0.504 0.754
QuiltNet [16] 0.403 0.318 0.491 0.564 0.564 0.591 0.884 0.801 0.578 0.435 0.322 0.467 0.364
MLLM4PUE-Q 0.437 0.434 0.868 0.851 0.686 0.668 0.886 0.866 0.613 0.516 0.440 0.507 0.758
E5-V [18] 0.251 0.324 0.367 0.169 0.299 0.306 0.500 0.348 0.385 0.166 0.337 0.053 0.267
CONCH [31] 0.559 0.444 0.832 0.802 0.703 0.644 0.890 0.665 0.439 0.447 0.405 0.607 0.783
PLIP [15] 0.251 0.306 0.708 0.253 0.532 0.541 0.721 0.749 0.551 0.488 0.179 0.290 0.450
MLLM4PUE-O 0.569 0.448 0.793 0.429 0.666 0.546 0.958 0.775 0.624 0.518 0.426 0.453 0.671
PathCLIP [37] 0.556 0.386 0.784 0.666 0.610 0.489 0.937 0.888 0.698 0.350 0.362 0.378 0.688
MLLM4PUE-P 0.599 0.386 0.812 0.892 0.767 0.784 0.937 0.886 0.744 0.576 0.509 0.496 0.760
QuiltNet [16] 0.441 0.413 0.628 0.503 0.607 0.586 0.904 0.802 0.588 0.434 0.338 0.458 0.449
MLLM4PUE-Q 0.479 0.441 0.866 0.849 0.695 0.661 0.887 0.872 0.609 0.494 0.444 0.520 0.761
Table 3: Performance comparison with baseline methods on zero-shot classification tasks, reporting weighted F1 (wF1) (top) and accuracy (bottom) metrics. MLLM4PUE-O, MLLM4PUE-P, and MLLM4PUE-Q denote our model pre-trained on Openpath [15], PathCap [37], Quilt1M [16], respectively. Bold values indicate the best performance. Same color represents models trained on the same dataset.